# Dynamical complexity of non-Gaussian many-body systems with dissipation

### v1

@@ Abstract
Introduction. Whether many-body quantum systems evolve into classically non-trivial states in the presence of a decohering environment is not only of fundamental interest to the theory of open quantum systems, but also has implications for the quantum advantage achievable by quantum computers and simulators [1, 2, 3, 4, 5, 6]. Traditionally, this question is often theoretically analyzed for many-body spin models which describe some experimentally accessible many-body systems very well. Here, there has been extensive recent activity in both the discrete-time setting (i.e. quantum circuits interspersed with noise events) and in the continuous-time setting (modeled by a many-body Lindblad master equation [7, 8]). For the discrete-time models, early results showed that a sufficiently high rate of noise prevents a buildup of entanglement in the quantum circuit and renders it classically simulable [9]. Recently, more refined results have been obtained showing classical simulability, in the presence of even a small amount of depolarizing noise, for either sampling or computing local observables in both random [10, 11, 12, 13, 14, 15, 16, 17, 18, 19] and non-random models [20, 21, 22]. Some of these results have also been generalized to the geometrically-local continuous-time setting, which more accurately models analog quantum simulators, and it has been established that the system remains classically simulable when the noise rate is higher than a threshold determined by the strength of the interaction terms in the Hamiltonian [8].
@@
Quantum simulators based on platforms such as ultracold atoms in optical lattices [23, 24, 25, 26, 27], superconducting circuits [28, 29] or nonlinear photonics [30, 31, 32], are often described by a family of Hamiltonians that, only in certain parameter regimes, reduce to quantum spin systems. The underlying physical model often comprises a lattice of fermionic or bosonic modes with two kinds of terms: (i) Gaussian coupling terms which are linear or quadratic in creation and annihilation operators and describe, for instance, particle hopping or pair production; (ii) Non-Gaussian interacting terms that typically act on particles only on one site. Consequently, there are two frequency scales in these Hamiltonians: the strength of the Gaussian coupling terms, $J$, and that of the on-site (non-Gaussian) interactions, $U$. When either $J=0$ or $U=0$, this model can be simulated classically. Specifically, when $J=0$, the model is a sum of single-site terms which map an initial product state to a product state from which we can both sample and compute local observables classically. When $U=0$, while time evolution can generate entangled states from an initial product state, these states remain Gaussian as long as the initial state is Gaussian. It is well-known that local observables can be efficiently computed in such states, and in the case of fermions, even sampling on the Fock state basis can also be efficiently performed [33, 34, 35, 36]. However, when both $J,U\neq 0$ and the system is noiseless, then it is well known that this model is universal for quantum computation [37, 38] and thus worst-case hard to simulate classically. While it is expected that the presence of dissipation should make the model classically easy to simulate, the amount and type of local dissipation needed remains less well understood.
@@
For fermionic systems, the impact of noise on the non-Gaussianity in the state was studied in a circuit model comprising only Gaussian gates and ancillas in a non-Gaussian resource state, and it was established that, independent of the gates and above a noise threshold, the state of the fermions is a convex combination of Gaussian states [39, 40, 41]. Studies analyzing continuous-time dynamics have been focused on problems with Gaussian Hamiltonians and the non-Gaussianity introduced via two-body dissipation [7]. For bosonic systems, previous studies have either focused on understanding their complexity classification as a function of evolution time in the absence of noise [42, 43, 44], or for the specific task of boson sampling in the presence of noise [45, 46, 47, 48, 49]. To the best of our knowledge, understanding classical simulability of noisy bosonic and fermionic continuous-time dynamics remains open.
@@
In this Letter, we rigorously address this question—we consider fermionic and bosonic systems with $n$ sites, each containing locally $L$ modes (Fig. 1). The annihilation operators corresponding to the $\sigma^{\text{th}}$ mode at the $i^{\text{th}}$ site, where $\sigma\in\{1,2\dots L\}$ and $i\in\{1,2\dots n\}$, is given by $a_{i,\sigma}$. It will be convenient to use the Hermitian operators $c_{i,\sigma}^{\alpha}$, with $\alpha\in\{1,2\}$, defined as $c_{i,\sigma}^{1}=(a_{i,\sigma}+a_{i,\sigma}^{\dagger})/\sqrt{2},c_{i,\sigma}^{2}=-i(a_{i,\sigma}-a_{i,\sigma}^{\dagger})/\sqrt{2}$, which represent either Majorana operators (for fermions) or position and momentum quadrature operators (for bosons). The noisy dynamics under consideration is described by the Lindblad master equation
$\frac{d\rho(t)}{dt}=-i[H(t),\rho(t)]+\kappa\sum_{i,\sigma}\mathcal{L}_{i,\sigma}\rho(t),$ (1)
@@
were $H(t)$ is a (possibly time-dependent) Hamiltonian that represents the unitary evolution of the system, and $\mathcal{L}_{i,\sigma}$ captures the noise, which is assumed to act locally on every mode $(i,\sigma)$ at a rate $\kappa$. We model the noise Lindbladian $\mathcal{L}_{i,\sigma}$ by
$\mathcal{L}_{i,\sigma}(\cdot)=\sum_{l=1}^{3}\xi^{(l)}\bigg{(}L_{i,\sigma}^{(l)}(\cdot)L^{(l)\dagger}_{i,\sigma}-\frac{1}{2}\{L_{i,\sigma}^{(l)\dagger}L_{i,\sigma}^{(l)},(\cdot)\}\bigg{)},$ (2)
@@
where we assume a physically-motivated and generic noise model with jump operators $L_{i,\sigma}^{(1)}=a_{i,\sigma},L_{i,\sigma}^{(2)}=a_{i,\sigma}^{\dagger},L_{i,\sigma}^{(3)}=a_{i,\sigma}^{\dagger}a_{i,\sigma}$ with relative decay rates $\xi^{(1)},\xi^{(2)},\xi^{(3)}$ respectively. The jump operator $a_{i,\sigma}$ models particle loss, $a_{i,\sigma}^{\dagger}$ models incoherent particle gain, and $a_{i,\sigma}^{\dagger}a_{i,\sigma}$ models dephasing. All of these noise processes occur in experiments, and it is often the case that $\xi^{(1)},\xi^{(2)},\xi^{(3)}>0$. While we choose this dissipation model for concreteness, our conclusions will continue to hold for other physically relevant dissipators such as $L_{i,\sigma}^{(1)}=c_{i,\sigma}^{1}$, $L_{i,\sigma}^{(2)}=c_{i,\sigma}^{2}$ which in the bosonic case would correspond to white noise fluctuations in the quadratures. We use the same noise Lindbladian for both bosons and fermions—for the bosonic case, we will additionally assume that the particle loss occurs at a rate strictly higher than (both coherent and incoherent) particle gain (see Supplemental Material [50] for the exact assumption) so as to avoid an unbounded growth of the number of particles with $t$ which would be unphysical in an actual experiment.
@@
and $H_{\text{ng}}(t)$ contains on-site non-Gaussian terms which account for particle-particle repulsion and attraction between different fermionic or bosonic modes at the same site:
@@
Note that in the fermionic case $\Omega_{i;\sigma}^{\alpha}(t)=0$, since physical Hamiltonians must preserve fermionic parity, while in the bosonic case $\Omega_{i;\sigma}^{\alpha}(t)$ can be a non-zero real scalar. For fermions, we can always assume that $J_{i,\sigma;j,\sigma^{\prime}}^{\alpha,\alpha^{\prime}}(t)$ is purely imaginary and anti-symmetric i.e.
@@
Figure 1: Sketch of the system that we study, with $n$ sites on a lattice, where each site contains $L$ modes. We represent a case with $L=5$ for simplicity. There are Gaussian couplings (in red) between the different sites, while the non-Gaussian interactions (in blue) are only onsite. Nonlocal couplings are allowed. The noise acting on each mode is represented by arrows. In the bosonic case, interactions of the form $n_{i,\sigma}^{2}$ are also allowed.
@@
We assume that the initial state $\rho(0)$ is either a product state (when analyzing the entanglement), a Gaussian state (when analyzing the non-Gaussianity), or both (such as the vacuum state). In the bosonic case, additionally, $\rho(0)$ will be assumed to satisfy $\mathrm{Tr}(n_{i,\sigma}^{k}\rho(0))\leq C_{0}^{k}k^{\alpha_{0}k+\beta_{0}},\forall(i,\sigma)$, for $k\in\{1,2,3...\}$, and for some $C_{0},\alpha_{0},\beta_{0}>0$: this condition guarantees that the probability of finding $\geq k$ particles in a mode decreases super-polynomially with $k$, as would be expected in a physically relevant bosonic state [51].
In the remainder of this paper, we analyze the dynamics of both bosonic and fermionic models and show that dissipation could make it easy to simulate this system when both $U,J\neq 0$ [Fig. 2]. First, we find that in the fermionic case, for a dephasing rate larger than the non-Gaussian interaction strength ($\kappa\xi^{(3)}\geq 2U$), the state remains convex-Gaussian at all times, irrespective of the value of $J$ or $\xi^{(1)},\xi^{(2)}$. Furthermore, we show that this convex-Gaussian state can be efficiently classically sampled from (in the Fock state basis). Furthermore, we also find that in the bosonic case, for particle loss and particle gain rates larger than the Gaussian coupling strength ($\kappa\geq 2J/\min(\xi^{(1)},\xi^{(2)})$), the state remains separable at all times regardless of the value of $U$ or $\xi^{(3)}$. Furthermore, there is an efficient classical algorithm that can allow us to sample (again, in the Fock state basis) from this separable state. We remark that our analysis provides a sufficient condition for simulability that applies to $any$ instance: it thus constitutes a worst-case analysis.
Our analysis also reveals fundamental differences between the high-noise behavior of bosonic and fermionic models. First, we show that unlike in the bosonic model, the fermionic model can evolve into entangled states at short-times even when the dissipation is much larger than the Gaussian couplings $\kappa\min(\xi^{(1)},\xi^{(2)},\xi^{(3)})\gg J$. Second, unlike the fermionic model, we provide evidence that the bosonic model can evolve into states that are not convex-Gaussian even with noise rates much larger than the non-Gaussian interactions $\kappa\min(\xi^{(1)},\xi^{(2)},\xi^{(3)})\gg U$, provided that the Gaussian couplings can be made sufficiently high.
Figure 2: Phase diagram for both bosonic and fermionic systems in the presence of generic noise. (a) For fermionic system the state remains convex Gaussian at all times for error rates $\kappa\xi^{(3)}\geq 2U$. (b) In bosonic systems the state remains separable at all times for error rates $\kappa\geq 2J/\min(\xi^{(1)},\xi^{(2)})$.
Results: Our results analyze the simulability of the fermionic and bosonic models in relation to the strength of the Gaussian couplings, non-Gaussian interactions, and noise rate. We first establish that when the noise rate is larger than the on-site non-Gaussian interaction strength $U$, the fermionic model remains convex-Gaussian at all times, and can therefore be classically efficiently sampled from.
Figure 3: Schematic depiction of the Trotterization schemes in the proof of Theorem 1 (fermionic systems with weak non-Gaussianity). For simplicity, we only depict a 1D setting, with each site containing $2$ modes ($L=2$). A single Trotter step consists of a Gaussian channel (red rectangles), followed by non-Gaussian gates (blue rectangles) interspersed with noise (gray circles). Crucially, a non-Gaussian gate followed by sufficiently strong noise can be written as a convex combination of Gaussian channels.
@@
For an initial Gaussian state, if $\kappa\geq 2U/\xi^{(3)}$, then the state of the fermionic model at time $t$, $\rho(t)$, is a convex combination of Gaussian states for all $t\geq 0$. Furthermore, $\rho(t)$ can be classically sampled in the Fock state basis to an $\epsilon$ total variation error in $\textnormal{poly}(n,t,1/\varepsilon)$ time.
Physically Theorem 1 suggests that if the noise rate is larger than the non-Gaussian interaction strength $U$, the non-Gaussianity is destroyed faster than it is created, and hence the classical simulation remains tractable. Since Gaussian terms in the Hamiltonian, no matter how strong, cannot make a convex-Gaussian state non-Gaussian, the noise threshold in Theorem 1 is independent of $J$ (the inter-site Gaussian interaction strength). Perhaps surprisingly, it is dephasing that results in convex Gaussianity at all times: at short times, incoherent particle loss and gain cannot preserve convex-Gaussianity in the presence of $H_{\text{ng}}(t)$ irrespective of $\kappa$. This is due to the fact that for the dynamics to be convex-Gaussianity preserving, it has to be convex-Gaussianity preserving separately in the even and odd parity subspaces [40]. However, while $H_{\text{ng}}(t)$ is parity conserving and acts individually in the two parity subspaces, the particle loss or gain dissipators are not parity number conserving and instead of acting individually in the two parity subspaces, couple the two parity subspaces. Consequently, these noise processes are unable to counter the non-convex-Gaussianity created by $H_{\text{ng}}(t)$ within the two parity subspaces at least at short times.
@@
We provide a complete proof of Theorem 1 in the supplement and only sketch the key ideas here. The starting point of showing Theorem 1 is a Trotterization of the Lindbladian in Eq. 1 — in each Trotter step, we express the evolution as (a) a Gaussian unitary corresponding to $H_{\text{g}}(t)$ as well as the particle loss and gain dissipators followed by (b) the single-site channels generated by $H_{\text{ng}}(t)$ together with the dephasing noise (Fig. 3). A close analysis of the single-site channel allows us to show that for $\kappa$ above the threshold in Theorem 1, this channel maps an input convex-Gaussian state to a convex-Gaussian state. Furthermore, we explicitly construct an approximation to this convex-Gaussian state, which allows us to use it to sample the state [35, 52].
@@
Next, we consider the bosonic model and establish that when the noise rate is larger than the inter-site Gaussian coupling, the bosonic model remains separable at all times, and can therefore be classically efficiently sampled from in the Fock state basis.
@@
Figure 4: Schematic depiction of the Trotterization schemes in the proof of Theorem 2 (bosonic systems with weak inter-site couplings). For simplicity, we only depict a 1D setting, with each site containing $2$ modes ($L=2$). A single Trotter step consists of a layer of single-mode channels, followed by 2-site gates interspersed with noise (in gray circles). Crucially, a 2-site gate followed by sufficiently strong noise can be written as a convex combination of single-site channels
@@
Suppose $\rho(t)$ is the state obtained after evolving the bosonic model for time $t$ with an initial product state, then for $\kappa\geq 2J/\min(\xi^{(1)},\xi^{(2)})$ the state $\rho(t)$ is separable for all $t\geq 0$. Furthermore, there is a randomized classical algorithm that can sample $\rho(t)$ in the Fock state basis to $\epsilon$ total variation error in $\textnormal{poly}(n,t,1/\epsilon)$ time.
Physically, our result makes precise the intuitive expectation that when the noise rate exceeds the inter-site interaction strength, the build up of entanglement is prohibited by the noise and thus, in this regime, a classically non-trivial entangled state is not generated. Notably, dephasing noise does not appear in the theorem. In fact, starting with a state which is diagonal in the Fock basis, one can easily check that entanglement can be generated for short times in the presence of depahsing noise, since dephasing noise leaves diagonal states invariant. Hence, dephasing alone cannot prevent entanglement buildup at all times. In the Supplemental Material, we also extend Theorem 2 to other experimentally relevant settings, such as for quantum spin models with single spin noise or for bosonic models with possible inter-site non-Gaussian couplings [50]. In contrast to previous results for spin systems which rely on percolation arguments and show that, at large noise rates, only $O(\log n)$ clusters of qubits can be entangled [8, 9], we instead show that the state is entirely separable i.e. can be written as a convex combination of product states at all sites. Furthermore, by providing an explicit construction of this separable state, we also show that this separable state can be efficiently sampled from.
@@
A detailed proof of Theorem 2 is provided in the Supplemental Material [50]. Similar to Theorem 1, we begin by a first-order Trotterization of the model but with a different decomposition of the Lindbladian: We express it as a product of (a) single site gates, which contain the unitary generated by the $H_{\text{ng}}(t)$ and the single-site terms in $H_{\text{g}}(t)$ and (b) two-site channels which contain the unitary generated by the inter-site couplings in $H_{\text{g}}(t)$ paired together with the noise [Fig. 4]. We denote the channel acting between modes $(i,\sigma)$ and $(j,\sigma^{\prime})$ at the Trotter time-step $\tau$ as $\Phi_{\tau\delta,(\tau-1)\delta}^{i,\sigma;j,\sigma^{\prime}}$, where $\delta$ is the size of the Trotter step. Importantly, the Trotterization is performed in such a way that the channel $\Phi_{\tau\delta,(\tau-1)\delta}^{i,\sigma;j,\sigma^{\prime}}$ can be understood as a time evolution of the inter-site Gaussian couplings between modes $(i,\sigma)$ and $(j,\sigma^{\prime})$, followed by noise on both modes. This choice of ordering the Trotterized evolution effectively redistributes the single-site noise into “gate-based” noise on the inter-site gates. Then, we explicitly analyze the two-site channel $\Phi_{\tau\delta,(\tau-1)\delta}^{i,\sigma;j,\sigma^{\prime}}$ and establish that for $\kappa$ above the threshold in Theorem 2, it maps a separable state into a separable state. Intuitively, a two-site gate followed by sufficiently strong noise loses its ability to generate entanglement. This allows us to show that the Trotterized state remains separable at all times for noise rates above a threshold.
Furthermore, in addition to certifying that the state is separable at all times, our analysis also allows us to explicitly construct an $O(\delta^{2})$ approximation to the separable state after each time step thus yielding an explicit algorithm to sample from the full quantum state. We remark that this separability does not come from the Trotterized noise channel becoming entanglement breaking above this threshold: since $\delta=\Theta(1/\text{poly}(n))$ for the Trotterization to be accurate, consequently $\Phi_{\tau\delta,(\tau-1)\delta}^{i,\sigma;j,\sigma^{\prime}}\to\textnormal{id}$ in the limit of large system-size (here id stands for identity). Our analysis thus has to explicitly capture the competition between the entanglement generated by the inter-site unitary and the simultaneous disentangling action of the noise.
@@
Tightness of Theorems 1 and 2. We can now ask if a version of Theorem 1 holds for bosonic systems i.e., is there a noise threshold $\kappa_{\text{th}}(U)$ dependent only on the non-Gaussian strength $U$ and uniform in $J,\Omega$, such that the state of the system becomes convex-Gaussian at all times? In the presence of dephasing noise $\xi^{(3)}\geq 0$, we provide numerical evidence to the contrary: with dephasing noise rates larger than the non-Gaussianity $(\kappa\xi^{(3)}\geq U)$, even a single bosonic mode can reach states which have negative Wigner functions and are thus not convex Gaussian [53, 50]. Therefore, as opposed to the fermionic case, dephasing noise in bosonic systems alone is not sufficient to remove the non-Gaussianity that is being introduced. While we cannot rule out the existence of a classical algorithm to simulate the system in this setting, in general, Wigner negativity is understood as a necessary resource for non-classicality and quantum advantage [53, 54], which suggests that classical simulation might be hard. In the absence of dephasing noise $(\xi^{(3)}=0)$, an even stronger argument can be made: in the Supplemental Material we argue that for any fixed $U$ no matter how small relative to $\kappa$, even the task of computing expectation values of local particle numbers is BQP-hard if we allow $J,\Omega$ to be arbitrarily large but $O(1)$ [50]. This builds upon Refs. [55, 56] which used large displacements to perform a universal set of gates on a single bosonic mode, such that the effective rate of noise on the implemented gate could be made arbitrarily small by increasing the displacement and carefully choosing a squeezed drive on the mode. We extend this technique to also implement an entangling gate between two oscillators, hence allowing for the implementation of fast, high-fidelity arbitrary gates. Together with results from Ref. [57], this suggests that when the noise is non-unital (i.e. $\kappa\xi^{(1)}\neq\kappa\xi^{(2)}$), by using sufficiently large $J$ and $\Omega$, a universal gate-set with effective gate noise less than the fault-tolerance threshold can be encoded into the Hamiltonian $H(t)$ [58, 59, 60]. This would imply that it is unlikely to be able to classically compute even local observables in the bosonic model unless BQP = BPP.

### v2

@@ Abstract
Introduction. Whether many-body quantum systems evolve into classically non-trivial states under decoherence is of fundamental interest to the theory of open quantum systems and also has implications for the quantum advantage in quantum computers and simulators [1, 2, 3, 4, 5, 6]. Traditionally, this has been mostly studied for many-body spin models, including extensive recent activity in both the discrete-time setting (i.e. quantum circuits interspersed with noise) and in the continuous-time setting (modeled by a many-body Lindblad master equation [7, 8]). For discrete-time models, early results showed that sufficiently high noise suppresses entanglement thus enabling classical simulation [9]. Recent results have shown classical simulability, even with a small amount of depolarizing noise, for both sampling or computing local observables in both random [10, 11, 12, 13, 14, 15, 16, 17, 18, 19] and structured models [20, 21, 22]. These results have partly been generalized to the geometrically-local continuous-time setting, which more accurately models analog quantum simulators, to show that the system remains classically simulable when the noise rate is larger than the interaction terms in the Hamiltonian [8].
@@
Quantum simulators based on platforms such as ultracold atoms in optical lattices [23, 24, 25, 26, 27], superconducting circuits [28, 29] or nonlinear photonics [30, 31, 32], are often described by a family of Hamiltonians that, only in certain regimes, reduce to spin systems. They are modeled by a fermionic or bosonic lattice with two kinds of terms: (i) Gaussian coupling terms which are linear or quadratic in creation/annihilation operators (e.g. particle hopping or pair production); (ii) Non-Gaussian interaction terms that typically act on particles only on one site. Consequently, there are two relevant frequency scales: the strength of the Gaussian couplings, $J$, and that of the on-site (non-Gaussian) interactions, $U$. If $J=0$ or $U=0$, this model is classically simulable. When $J=0$, the Hamiltonian is a sum of single-site terms which maps product states to product states which can be classically simulated. When $U=0$, the dynamics is Gaussian and thus local observables can be efficiently computed and, for fermions, even sampling is efficient [33, 34, 35, 36]. However, when both $J,U\neq 0$ and the system is noiseless, this model is universal for quantum computation [37, 38] and thus worst-case hard to simulate classically. While the presence of dissipation should make the model classically simulable, the amount and type of local dissipation needed remains unclear.
@@
For fermionic systems, the impact of noise on non-Gaussianity was studied in a circuit model with Gaussian gates and non-Gaussian ancillas, which showed that the state remains a convex combination of Gaussian states above a noise-threshold[39, 40, 41]. Studies analyzing continuous-time dynamics have focused on the non-Gaussianity introduced via two-body dissipation [7]. For bosonic systems, previous studies have either focused on understanding their complexity as a function of evolution time in the absence of noise [42, 43, 44], or for the specific task of boson sampling in the presence of noise [45, 46, 47, 48, 49]. However, the classical simulability of the noisy continuous-time model motivated above remains unresolved.
@@
In this Letter, we rigorously address this question—we consider fermionic and bosonic systems with $n$ sites, each containing locally $L$ modes (Fig. 1). The annihilation operators corresponding to the $\sigma^{\text{th}}$ mode at the $i^{\text{th}}$ site, where $\sigma\in\{1,2\dots L\}$ and $i\in\{1,2\dots n\}$, is given by $a_{i,\sigma}$. We will use the Hermitian operators $c_{i,\sigma}^{\alpha}$, with $\alpha\in\{1,2\}$, defined as $c_{i,\sigma}^{1}=(a_{i,\sigma}+a_{i,\sigma}^{\dagger})/\sqrt{2},c_{i,\sigma}^{2}=-i(a_{i,\sigma}-a_{i,\sigma}^{\dagger})/\sqrt{2}$, which represent either Majorana operators (for fermions) or position and momentum quadrature operators (for bosons). The noisy dynamics is described by the Lindblad master equation
$\frac{d\rho(t)}{dt}=-i[H(t),\rho(t)]+\sum_{i,\sigma}\mathcal{L}_{i,\sigma}\rho(t),$ (1)
@@
were $H(t)$ is a (possibly time-dependent) Hamiltonian describing the system. $\mathcal{L}_{i,\sigma}$ captures the noise, which is assumed to act locally on every mode $(i,\sigma)$, and is modeled by
$\mathcal{L}_{i,\sigma}(\cdot)=\sum_{l=1}^{3}\kappa_{l}\bigg(L_{i,\sigma}^{(l)}(\cdot)L^{(l)\dagger}_{i,\sigma}-\frac{1}{2}\{L_{i,\sigma}^{(l)\dagger}L_{i,\sigma}^{(l)},(\cdot)\}\bigg),$ (2)
@@
with jump operators $L_{i,\sigma}^{(1)}=a_{i,\sigma},L_{i,\sigma}^{(2)}=a_{i,\sigma}^{\dagger},L_{i,\sigma}^{(3)}=a_{i,\sigma}^{\dagger}a_{i,\sigma}$ and decay rates $\kappa_{1},\kappa_{2},\kappa_{3}$ respectively. The jump operator $a_{i,\sigma}$ models particle loss, $a_{i,\sigma}^{\dagger}$ models incoherent particle gain, and $a_{i,\sigma}^{\dagger}a_{i,\sigma}$ models dephasing. We remark that our conclusions also hold for other physically relevant dissipators such as $L_{i,\sigma}^{(1)}=c_{i,\sigma}^{1}$, $L_{i,\sigma}^{(2)}=c_{i,\sigma}^{2}$ which, in the bosonic case, would correspond to white noise fluctuations in the quadratures. We use the same noise Lindbladian for both bosons and fermions—for the bosonic case, we will additionally assume that the particle loss occurs at a rate strictly higher than (both coherent and incoherent) particle gain (see Supplement [50] for the exact assumption) so as to avoid an unbounded growth of the number of particles with $t$ which would be unphysical in an actual experiment.
@@
and $H_{\text{ng}}(t)$ contains on-site non-Gaussian terms which account for particle-particle repulsion and attraction between different modes at the same site:
@@
Note that in the fermionic case $\Omega_{i;\sigma}^{\alpha}(t)=0$, since physical Hamiltonians must preserve fermionic parity, while in the bosonic case $\Omega_{i;\sigma}^{\alpha}(t)$ can be a non-zero real scalar. For fermions, we can assume that $J_{i,\sigma;j,\sigma^{\prime}}^{\alpha,\alpha^{\prime}}(t)$ is purely imaginary and anti-symmetric i.e.
@@
Figure 1: Sketch of the model, with $n$ sites on a lattice, where each site contains $L$ modes. There are Gaussian couplings between the different sites, while the non-Gaussian interactions are only onsite. Nonlocal couplings are allowed. In the bosonic case, interactions of the form $n_{i,\sigma}^{2}$ are also allowed.
@@
The initial state $\rho(0)$ is either a product state (when analyzing entanglement), a Gaussian state (when analyzing non-Gaussianity), or both (e.g., the vacuum state). In the bosonic case, additionally, $\rho(0)$ will be assumed to satisfy $\mathrm{Tr}(n_{i,\sigma}^{k}\rho(0))\leq C_{0}^{k}k^{\alpha_{0}k+\beta_{0}},\forall(i,\sigma)$, for $k\in\{1,2$...\}, and for some $C_{0},\alpha_{0},\beta_{0}>0$: this condition guarantees that the probability of finding $\geq k$ particles in a mode decreases super-polynomially with $k$, as would be expected in a physically preparable bosonic state [51].
Figure 2: Phase diagram for both bosonic and fermionic systems in the presence of generic noise. (a) For fermionic system the state remains convex-Gaussian at all times for error rates $\kappa_{3}\geq 2U$. (b) In bosonic systems the state remains separable at all times for error rates $\kappa_{1},\kappa_{2}\geq 2J$.
Results: Our results, depicted in Fig. 2, show the simulability of the fermionic and bosonic models as a function of $J,U,\kappa_{i}$. We first establish that when the noise rate is larger than the on-site non-Gaussian interaction strength $U$, the fermionic model remains convex-Gaussian at all times, and can therefore be classically efficiently sampled from.
Figure 3: Schematic depiction of the Trotterization schemes in the proof of Theorem 1 (fermionic systems with weak non-Gaussianity). For simplicity, we only depict a 1D setting, with each site containing 33 modes ($L=3$). A single Trotter step consists of a Gaussian channel (orange rectangles) that includes the combined effect of $H_{g}(t)$ and the particle gain and loss dissipators, followed by non-Gaussian gates (blue rectangles) interspersed with the dephasing dissipator (gray curved rectangles). Crucially, a non-Gaussian gate followed by sufficiently strong dephasing can be written as a convex combination of Gaussian channels.
@@
For an initial Gaussian state, if $\kappa_{3}\geq 2U$, then the state of the fermionic model at time $t$, $\rho(t)$, is a convex combination of Gaussian states for all $t\geq 0$. Furthermore, $\rho(t)$ can be classically sampled in the Fock state basis to an $\epsilon$ total variation error in $\textnormal{poly}(n,t,1/\epsilon)$ time.
Physically Theorem 1 suggests that when $\kappa_{3}\geq 2U$, dephasing noise destroys non-Gaussianity faster than $H_{\text{ng}}(t)$ creates it, so that $\rho(t)$ always remains convex-Gaussian. Since $H_{\mathrm{g}}(t)$ preserves convex-Gaussianity, the noise threshold in Theorem 1 is independent of $J$. Notably, it is dephasing that results in this convex-Gaussianity. With only incoherent particle loss/gain, $\rho(t)$ could evolve into a non-Gaussian state at short times even with large dissipation. This arises from the fermionic parity structure—the density matrix of the fermionic model has the form $\rho(t)=\rho_{+}(t)+\rho_{-}(t)$, where $\rho_{\pm}(t)$ is supported only on even/odd parity states. Due to this structure, convex-Gaussianity in $\rho(t)$ requires both $\rho_{\pm}(t)$ to be convex-Gaussian [40]. $H_{\text{ng}}(t)$ generates non-Gaussianity individually in both $\rho_{\pm}(t)$. However, the loss/gain dissipators, to first order, switch the parity of the state and do not act within the two parity subspaces. Consequently, they cannot immediately counter the non-convex-Gaussianity created by $H_{\text{ng}}(t)$.
@@
We provide a complete proof of Theorem 1 in the supplement: The starting point is a Trotterization of the Lindbladian in Eq. 1 — in each Trotter step, we express the evolution as (a) a Gaussian unitary corresponding to $H_{\text{g}}(t)$, particle loss and gain followed by (b) the single-site channels generated by $H_{\text{ng}}(t)$ and dephasing (Fig. 3). Analyzing the single-site channel, we show that for $\kappa_{3}\geq 2U$, this channel maps an input convex-Gaussian state to a convex-Gaussian state. Furthermore, we explicitly construct the output convex-Gaussian state, which allows us to sample from it [35, 52].
@@
Next, we consider the bosonic model and establish that when the noise rate is larger than the inter-site Gaussian coupling, $\rho(t)$ remains separable at all times, and can therefore be classically efficiently sampled from in the Fock state basis.
@@
Figure 4: Schematic depiction of the Trotterization schemes in the proof of Theorem 2 (bosonic systems with weak inter-site couplings). For simplicity, we only depict a 1D setting, with each site containing 33 modes ($L=3$). A single Trotter step consists of a layer of single-site channels which include the Hamiltonian terms acting on that site and the dephasing dissipator, followed by 2-site gates interspersed with particle gain and loss dissipators (in gray circles). Crucially, a 2-site gate followed by sufficiently strong noise can be written as a convex combination of single-site channels
@@
Suppose $\rho(t)$ is the state obtained after evolving the bosonic model for time $t$ with an initial product state, then for $\kappa_{1},\kappa_{2}\geq 2J$ the state $\rho(t)$ is separable for all $t\geq 0$. Furthermore, there is a randomized classical algorithm that can sample $\rho(t)$ in the Fock state basis to $\epsilon$ total variation error in $\textnormal{poly}(n,t,1/\epsilon)$ time.
Our result formalizes the intuition that, when noise exceeds the inter-site coupling strength, a buildup of entanglement is prohibited and no classically non-trivial state is generated. Notably, the noise threshold is determined by particle loss and gain noise. This arises from the dephasing dissipator being diagonal in the Fock basis, while evolution under $H_{g}(t)$ creates entanglement through off-diagonal elements (i.e., coherences). Thus, to first order, dephasing cannot counter this entanglement generation. Particle gain or loss dissipators, however, are not diagonal in the Fock basis and can prevent it. In the Supplement, we also extend Theorem 2 to quantum spin models with single spin noise and bosonic models with inter-site non-Gaussian couplings [50]. Unlike previous percolation-based arguments limiting entanglement to $O(\log n)$ qubit clusters for sufficiently strong noise [8, 9], we show that the state is entirely separable, and provide an explicit construction that can be efficiently sampled from.
@@
A detailed proof of Theorem 2 is provided in the Supplement [50]. Similar to Theorem 1, we begin by a first-order Trotterization of the model but with a different decomposition of the Lindbladian: We express it as a product of (a) single site gates, which contain the unitary generated by $H_{\text{ng}}(t)$, single-site terms in $H_{\text{g}}(t)$ and the dephasing dissipator and (b) two-site channels which contain the unitary generated by the inter-site terms in $H_{\text{g}}(t)$ paired together with the particle gain and loss dissipators [Fig. 4]. We denote the channel acting between modes $(i,\sigma)$ and $(j,\sigma^{\prime})$ at the Trotter-step $\tau$ as $\Phi_{\tau\delta,(\tau-1)\delta}^{i,\sigma;j,\sigma^{\prime}}$, where $\delta$ is the size of the Trotter-step. Importantly, the Trotterization is performed in such a way that the channel $\Phi_{\tau\delta,(\tau-1)\delta}^{i,\sigma;j,\sigma^{\prime}}$ can be understood as a time evolution of the inter-site Gaussian couplings between modes $(i,\sigma)$ and $(j,\sigma^{\prime})$, followed by noise on both modes. This effectively redistributes the single-site noise into “gate-based” noise on the inter-site gates. Analyzing $\Phi_{\tau\delta,(\tau-1)\delta}^{i,\sigma;j,\sigma^{\prime}}$, we show that for $\kappa_{1},\kappa_{2}\geq 2J$, it is separability-preserving and thus the state remains separable at all times. Furthermore, we also explicitly construct an $O(\delta^{2})$ approximation to the separable state after each time-step and obtain an explicit algorithm to sample from $\rho(t)$.
@@
Tightness of Theorems 1 and 2. We can now ask if a version of Theorem 1 holds for bosonic systems i.e., is there a noise threshold $\kappa_{\text{th}}(U)$ dependent only on the non-Gaussian strength $U$ and uniform in $J,\Omega$ that guarantees a convex-Gaussian at all times? For dephasing noise, we provide numerical evidence to the contrary: even for $\kappa_{3}\gg U$, single-mode dynamics can yield states with negative Wigner function, and thus not convex-Gaussian states can be generated at time-scales $\sim 1/U$ [53, 50]. While negativity of the Wigner function suggests classical simulation hardness [53, 54], we do not rule out the existence of an efficient classical algorithm. When $\kappa_{3}=0$, in the Supplement we show that no matter how small $U$ is relative to $\kappa_{1},\kappa_{2}$, computing expected local particle numbers is BQP-hard if $J,\Omega$ can be arbitrarily large but $O(1)$ [50]. This builds upon Refs. [55, 56] which perform a universal gate-set on a single bosonic mode with an arbitrarily small effective gate-error rate by engineering the displacement and squeezing. We extend this technique to also implement an entangling gate between two oscillators thus yielding a high-fidelity universal multi-mode gate-set. Together with results from Ref. [57], this suggests that when the noise is non-unital (i.e. $\kappa_{1}\neq\kappa_{2}$), by using sufficiently large $J$ and $\Omega$, a fault-tolerant quantum computation can be encoded into the model [58, 59, 60]. Thus, it is unlikely to be able to classically compute even local observables in this setting unless BQP = BPP.